OpenStudy (anonymous):

A bag contains 4 white, 3 blue, and 5 red marbles. Find the probability of choosing a blue marble, then a red marble if the marbles are replaced. A. 5/44 B. 15/35 C. 2/3 D. 1/15

OpenStudy (anonymous):

How many blue marbles are there? How many total marbles are there? And how many red marbles are there?

OpenStudy (anonymous):

3 blue, 12 total, and 5 red.

OpenStudy (anonymous):

So do \[\frac{ blue }{ total }*\frac{ red }{ total }\]

OpenStudy (anonymous):

15/144

OpenStudy (anonymous):

so it is B

OpenStudy (anonymous):

Well you can simplfy it to 5/48

OpenStudy (anonymous):

but there is no 5/48 on the choices. And how do you simplify?

OpenStudy (anonymous):

Are you sure the question doesn't say that they aren't replaced? Because we have the correct answer right now.

OpenStudy (anonymous):

im sure

OpenStudy (anonymous):

oh it says are not replaced, my bad

OpenStudy (anonymous):

Well then let's go again\[\frac{ blue }{ total }*\frac{ red }{ total-1 }\]

OpenStudy (anonymous):

15/143?

OpenStudy (anonymous):

\[\frac{ 3 }{ 12 }*\frac{ 5 }{ 11 }\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

15/132

OpenStudy (anonymous):

Now do\[\frac{ 15 }{ 132 }\div \frac{ 3 }{ 3 }\]

OpenStudy (anonymous):

5/44

OpenStudy (anonymous):

Boom, that's your answer

OpenStudy (anonymous):

thank you sooooo much!!!!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

kk bye